OpenStudy (anonymous):

quantity 8 x minus 56 over quantity x squared minus 49 divided by quantity x minus 6 over quantity x squared plus 11 x plus 28

OpenStudy (anonymous):

help?

OpenStudy (debbieg):

\[\frac{ (8x-56) }{x^{2}-49 }\div\frac{(x-6)}{x ^{2}+11x+28}\] =\[\frac{ 8(x-7) }{(x-7)(x+7)}\times\frac{(x+7)(x+4)}{(x-6)}\] =\[\frac{8(x+4)}{(x-6)}\]